OpenStudy (vortish):

-4(T -6) - (T+6)= 7 -4T -24 -T -6 = 7 - 5T - 30 = 7 + 30 + 30 - 5T = 37 is this going the right way pls help

OpenStudy (anonymous):

you made an error in line 2... you multiplied -4 by -6 in line 1 and replaced it with -24 in line 2... should have been +24

OpenStudy (compassionate):

-4t + 24 + t + 6 = 7 -4t + t+ 24 + 6 = 7 -3t + 30 = 7 -3t - 30 + 30 = 7 - 30 -3t = -23 Divide t = 7.8 Now switch t = 11/5

OpenStudy (anonymous):

fix that and then do the simplifying again...

OpenStudy (anonymous):

oops, @Compassionate is on it... thanks!

OpenStudy (anonymous):

Except that last step isn't right, I don't think... -3t = -23 divide both sides by -3 -3t/(-3) = -23/(-3) t = 23/3 That doesn't reduce any further... you can get it as a decimal if you want... 7.666667, but 23/3 is correct too.

OpenStudy (compassionate):

You can reduce 7.6 a little further to make it clear. But that's a different form of mathematics. I'd just go with 7.6

OpenStudy (anonymous):

I'm not sure I follow... 23/3 is equal to 7 and 2/3 as a mixed fraction, or you can write the 2/3 as a decimal, rounding off wherever you feel like. But you wouldn't want to round to 7.6. Anyway, the algebra is the more important part... @vortish, the steps you followed were good, and if you had not made that mistake early on, you probably would have been fine through the rest.

OpenStudy (compassionate):

That's very true, Jake, he only one mistake. ^ Look at me rhyming.